OpenStudy (anonymous):

Solve using the elimination method: 3x+4y=2 6x+8y=4

OpenStudy (anonymous):

multiply first equation by -2 and then add \[-6x-8x=-4\] \[6x+8x=4\] when you add you get \[0=0\] because everything drops out. that means they are the same line

OpenStudy (anonymous):

it does not mean there is no solution. nor does it mean that you cannot solve it. it means they are the same line, they just look different

OpenStudy (anonymous):

so any point on one line is also on the other. there are infinitely many solutions

OpenStudy (anonymous):

Satellite is right

OpenStudy (anonymous):

in fact before you do anything you see that there is no difference between \[3x+4y=2\] and \[6x+8y=4\] one is just twice the other on both sides, so any pair that fits one will fit the other

OpenStudy (anonymous):

I knew there was infinite solutions, I just didn't think it was solvable using the elimination method

OpenStudy (anonymous):

To reliably solve linear systems of equations you should always use row reduction (gaussian elimination)

OpenStudy (anonymous):

In this case you can just rearrange the first equation in terms of x and let y be arbitrary. But you would get this result with row reduction.

OpenStudy (anonymous):

x = (-4/3)y + 2/3

OpenStudy (anonymous):

Where y is any element of the field.